Left Direct Gregorian Technical Observing with IRTC: UT 2009 December 28

Observer: JHill (LBTO), DMiller (Tucson)
Telescope Operator: DHuerta
Instrument Support:
Telescope Support: BCurtis
SW support: (Tucson)
Phone Support:

Summary

First 2.5 hours to Andrew for Laser-Tracker measurements of collimation as a function of elevation.

Opened dome briefly, but then closed for 2.5 hours to fix hardpoint 2. See Issue #2448.

Reopened for 1 hour and were able to confirm that guide camera is now working in AGW4.

Closed for weather at 7:34 UT.

Preparations

Dan C. gave us a DSP configuration file that can run both rotators. But it still has problems with missing telemetry information that messes up the MCS rotator display. So we've switched back to the the LFBG-only configuration file to make things more clear.

Brian C. inspected AGW4, but didn't find any "obvious problems".

Telescope was parked at AZ=60 deg. Andrew says it looks spectacular from the Willcox side at sunrise.

The polycom unit has video again. (It was turned off for a few hours.)

Details

Laser Tracker data

00:20 Andrew is taking Laser-Tracker measurements as a function of elevation.

The telescope is in the best imaginable thermal equilibrium state for these measurements as it was at -6 degC all last night while we were observing, and the outside temperature only warmed slightly today. Telescope steel was at -3.5 degC during these measurements with a total gradient of about 1.3 degC from the C-ring to the windbracing. Things are in such good equilibrium with ambient that we decided it was pointless to try to take thermal gradient data as the dome opened.

03:00 Andrew is finished with data collection and dismounting his cables.

WFSC Tests

01:18 Took a LFBG WFSC image with IDL to confirm that there is enough light in the closed chamber to get a signal on the camera. Data in /OldRepository/AGW_Data/091228_wfs. LFBG image is wfscimage00001.fits.

01:25 Switch GCS to LDG with the /lbt/tcs/current/bin/setNewInstrument command
% setNewInstrument IRTC directGregorian
Return failure. My arguments must not be correct.

01:30 Tried to connect with vncviewer to agw4-azcamgw. failed. Logged into lbtmu107 as agwuser and did a startAGW -u 4. Could now connect.

01:44 sent setNewInstrument IRTC directGregorian Success

01:45 Took a LDG WFSC image with IDL and the image looks correct (could see circles in subapertures)

02:10 Took LDG Guider image with readGuideCam -e 2000 (and -e 1000) (dir above guider000001,2.fits) and it appears that the guide camera is working again. We see the dust particles that we saw when we took flats a month ago. So, I think we are going to try LDG Commissioning again.

Return to Direct

02:15 We got Dan C. to put us back on the DSP configuration that runs both rotators. This means that the telemetry status bits that are displayed by MCS are unreliable. MCSPU was restarted for the changes to take effect.

02:30 Changed GCS back to V2.2.8B4 as we are going to LDG.

02:35 Changed authorization back to IRTC at directGregorian.

02:58 Retracting the left M3 swing arm for Direct observations and rebalancing

Open

03:15 Open shutter doors T=-5 degC, D=-15 degC, wind 2-5 m/sec from WNW. Sky is covered with thin-medium cirrus, making for a spectacular 22-degree halo around the moon.

03:22 reset encoders at AZ=120 degrees.

IIF GUI Problem

The problem Doug saw with the IIF not authorizing the instrument and focal station that were displayed on the pull-down menus on the IIF GUI in Arcetri (AI LBTO #9) was seen again on the Mountain. For the following three "Authorize" requests from the IIF GUI Control window, the Left Instrument boxes displayed "IRTC" and "Direct". However, the first Authorize set the Instrument to LUCIFER and FS to bentFrontGregorian (seen in /lbt/log/2009/12/28 with output from PCS and GCS and in the event log with output from IIF.

From the event log:

03:24:16 Lucifer LFBG started (Success)

03:25:10 IRTC LDG started (Failed because rotator was not ready)

03:25:51 IRTC LDG started (Success)

Created IssueTrak #2469

First PresetTelescope

03:33 pointto M5_0162 TelMode=TRACK

03:36 PresetTelescope failed: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance. John finds that Hardpoint 2 is not moving.

Close

03:45 John and Brian make a visit to the SX Primary Cell (telescope at zenith). The hardpoint 2 stepper motor shaft was turning, but the hardpoint length was not changing even a little bit.

04:08 Apparently John and Brian are having a long heart-to-heart conversation with the primary cell.

06:30 John and Brian "fix" Hardpoint #2 via IssueTrak #2448

Open

06:37 pointto M5_0269 TelMode=TRACK gs=0

06:38 No star on IRTC. ptspiral one step. IE=21, CA=28

06:40 On-axis M5 star too bright

06:42 Several off-axis stars too dim.

06:46 pointto M5_0269 TelMode=Active gs=5. Success guiding and active optics. 14 iterations down to ~600 rms WFE. Seeing is from 1" to 1.5"

07:02 pointto M5_0347 TelMode=TRACK gs=0 Failed: Primary Mirror collimation reached a limit. David move primary in Y by changing the global offset from -2.633 to -2.000. (This was a deliberate non-standard procedure as we did not make the parallel move of M2, so the pointing was changed.)

07:05 pointto M5_0347 TelMode=TRACK gs=0 Success. Star on edge of IRTC. Set IE=41, CA=29 at EL=39

07:18 pointto BS9122 Star on side of Acq image (acquire000006.fits) but guided on noise instead. Torsten should check this image

07:25 pointto BS9122 TelMode=ACQUIRE Put star in center with IE=28, CA=20

07:28 pointto BS9122 TelMode=ACTIVE gs=0 Collimate wfsc 22-31. At wfsc 25, cleared Active Optic buffer on PSF GUI to restart colllimate. John thought that we may have had a "funny" M2 offload correction earlier. Collimation proceeded to 800nm RMS.

Close

07:34 Clouds getting thicker, a few snowflakes were seen so closed the chamber. T=-3 degC

-- JohnHill - 28 Dec 2009
Topic attachments
I Attachment Action Size Date Who Comment
daq.20091228.jpgjpg daq.20091228.jpg manage 645 K 29 Dec 2009 - 07:40 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.dx.thermal.20091228.jpgjpg pmcc.dx.thermal.20091228.jpg manage 522 K 29 Dec 2009 - 07:42 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20091228.jpgjpg pmcc.sx.thermal.20091228.jpg manage 555 K 29 Dec 2009 - 07:41 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r8 - 29 Dec 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback